javonarnold:

Simplify 15 to the eighteenth power over 15 to the third power . (4 points) 156 1515 16 115

Astrid1:

The simplified form is 15^{15} Step-by-step explanation: Consider the provided expression. 15 to the 18th power over 15 to the 3rd power. This can be written as: \frac{15^{18}}{15^{3}} Now use the rule of exponent: \frac{a^m}{b^n}=a^{m-n} By using the above rule of exponent. \frac{15^{18}}{15^3}=15^{18-3} \frac{15^{18}}{15^3}=15^{15} Hence, the simplified form is 15^{15} Site: https://istudy-helper.com/mathematics/question8061677#otvet

Kyt:

most versoza wants to compare simple and compound interest on a 3,500 pesos investment for three years find the interest if one's current 6.5% simple interest annually find the interest if found in 6.5% interest compounded annually find the difference between the two interest